If two fair dice are rolled, find the probability of the sum is 10 or both dice are odd.
Answer by stanbon(75887)

If two fair dice are rolled, find the probability of the sum is 10 or both dice are odd.
----------
Sketch the sample space as a 6 by 6 square
listing the 36 sums from 1+1=2 to 6+6=12.
Count the number of sum = 10
Count the number of both dice are odd
---
Remember P(A or B) = P(A)+P(B)-P(A and B)
=======================
Cheers,
Stan H.